our firm needs a computerized machine tool lathe which costs $78,463 and requires $12,287 in maintenance expense for each year of its 5-year life. After five years, this machine will be replaced. The machine falls into the MACRS (Links to an external site.) 5-year class life category. Assume a tax rate of 25 percent and a discount rate of 12 percent. If the lathe can be sold for $5,687 at the end of year 5, what is the after-tax salvage value? Below are the MACRS rates for the 5-year recovery period. DO NOT USE DOLLAR SIGNS OR COMMAS IN YOUR ANSWER. ENTER YOUR ANSWER TO THE NEAREST DOLLAR (e.g. 1250).
Answer #1

rate positively ..

Marcs depreciation rate table =
Year 5-Year
1 20.00%
2 32.00%
3 19.20%
4 11.52%
5 11.52%
6 5.76%
Book value at the end of 5th year= 78,463*5.76%
i Book value        4,519
ii Sales price 5687
iii=ii-i Profit on sale =        1,168
iv=iii*25% Tax on profit =            292
v=ii-iv After tax salvage value =        5,395